PROBLEM 2. A pyramidal monument has an octagonal base, each side of which is 2 m. Each lateral edge of the pyramid is 3m. Find the total cost of painting the lateral surface of the monument at PhP 730.00 per square meter?
FINAL ANSWER: _______________________________________
1
Expert's answer
2020-12-15T17:31:23-0500

"S=8\cdot S_1" is the area of the lateral surface of the monument

"S_1" is the area of one side of the monument

"P_1=730\space PhP" is the price per square meter

"P=S\cdot P_1" is the total cost of painting the lateral surface of the monument

"S_1=\sqrt{p\cdot(p-a)\cdot(p-b)\cdot(p-c)}"

"a=b=3"m

"c=2m"

"p=\frac {a+b+c}2=\frac {3+3+2}2=4m"

"S_1=\sqrt{p\cdot(p-a)\cdot(p-b)\cdot(p-c)}=\sqrt{4\cdot(4-3)\cdot(4-3)\cdot(4-2)}=2\sqrt {2}m^2"

"S=16\sqrt{2}m^2"

"P=S\cdot P_1=16\sqrt{2}\cdot730\approx16518\space PhP"